# MMDocIR: Benchmarking Multimodal Retrieval for Long Documents

### v2

@@ Abstract
$\dagger$\$$\dagger$\$footnotetext: These authors contributed equally
Figure 2: Area ratio of different modalities (1) in overall and (2) by domains in MMLongBench-Doc benchmark.
@@ 1 Introduction
Wiki-SS Natural question 3,610 TXT Wikipedia 1.0
@@
Multimodal document retrieval Hassan et al. (2013); Lee et al. (2024) aims to retrieve information from visually rich documents based on user queries. Unlike traditional document retrieval Zhang et al. (2022); Chen et al. (2023) which primarily deals with textual data, multimodal document retrieval imposes substantially greater demands on understanding multimodal elements such as images, tables, charts, and layout designs. Such elements often carry significant information that plain text fails to convey Cui et al. (2021); Sassioui et al. (2023): tables reveal structured data patterns, charts visualize trends or correlations, images offer contextual and semantic cues, etc. Combining these visual elements enriches the quality of retrieved content. Our analysis of MMLongBench-Doc benchmark Ma et al. (2024b) in Figure 2 shows that: text occupies only 52.7% of content area, while images and tables account for 29.2% and 12.8% respectively. This highlights the need for retrieval systems that effectively handle multimodal information.
@@
However, as shown on (see Table 1), existing benchmarks exhibit several critical limitations that undermine comprehensive evaluation of multimodal retrieval systems. The key limitations include: 1. Question Quality: Many questions used in existing benchmarks are directly sourced from datasets for Visual Question Answering (VQA) tasks. Some questions often assume the input is already relevant, making it not suited for meaningful evaluation of retrieval capabilities. 2. Document Completeness and Diversity: Existing benchmarks often provide only partial documents, limiting the ability to evaluate within full document context. Additionally, the narrow range of document domains further restricts their applicability across diverse use-cases in real-world. 3. Retrieval Granularity: Most benchmarks support only page-level retrieval. Such granularity is often insufficient, as user queries frequently target specific elements, such as figures or tables, rather than entire pages.
@@
To address these gaps, we introduce MMDocIR, a multimodal document information retrieval benchmark. MMDocIR is designed for two key tasks: page-level and layout-level retrieval. (1) The page-level retrieval identifies the most relevant pages within a document to answer user query. (2) The layout-level retrieval targets the most relevant fine-grained elements such as paragraphs, equations, figures, tables, and charts (see Appendix E.2 for more examples). Such task supports more precise and context-aware retrieval that pinpoint specific elements to address user queries. To support both tasks, we develop MMDocIR evaluation set that comprises 313 documents, each averaging 65.1 pages, along with 1,658 modified queries derived from MMLongBench-Doc and DocBench Zou et al. (2024). The queries are annotated with 2,107 page-level and 2,638 layout-level labels. The page labels are specific pages that contain the evidence needed to answer the query. [^1] While MMLongBench-Doc provided initial page labels, our meticulous review lead to corrections in 21.3% of them. The layout labels consist of precisely drawn bounding boxes around the key evidence within the identified pages. In addition, we introduce the MMDocIR training set, designed to support retriever training. It contains 73,843 questions sourced from 7 DocQA datasets. To construct this set, we manually collect 6,878 documents and apply a semi-automatic pipeline to annotate the ground truth labels.
@@
  • Dual-task Retrieval Framework: We propose a dual-task retrieval framework (§$\S$ 2) that supports page-level and fine-grained layout-level multimodal document retrieval.
  • MMDocIR Benchmark: We introduce Multimodal Document Information Retrieval benchmark. The evaluation set (§$\S$ 3) consists of 313 documents with expert-annotated labels for 1,658 questions. The training set (§$\S$ 4) consists of 6,878 documents and labels for 73,843 questions.
  • We conduct extensive experiments and comparisons of both text and visual retrievers (§$\S$ 5), demonstrating clear advantage of incorporating visual content in multimodal retrieval tasks.
@@ 2 Dual-Task Retrieval Definition
Let $\mathcal{D}$ be a document corpora consisting of document pages: $\mathcal{P}=\{p_{1},p_{2},\dots,p_{n}\}$, and layouts: =$\mathcal{L}=\{l_{1},l_{2},\dots,l_{m}\}$ extracted via layout detection. The objective is to perform document retrieval at both page-level and layout-level. Specifically, given query $Q$, the task is to retrieve the top $k$ pages and layouts most relevant to $Q$, where $k<<n$ and $k<<m$. The relevance of pages ($p$) and layouts ($l$) to $Q$ is measured by similarity scores, $\textrm{Sim}(Q,p)$ and $\textrm{Sim}(Q,l)$ respectively. The retrieval system consists of two phases: (1) an offline indexing phase, where pages and layouts from $\mathcal{P}$ and $\mathcal{L}$ are encoded into vectors, and (2) an online querying phase, in which a query $Q$ is encoded into a vector, which is then compared against the offline-indexed vectors using similarity scores $\textrm{Sim}(Q,p)$ for pages and $\textrm{Sim}(Q,l)$ for layouts.
@@ 3.2 Annotation Process
We annotate page labels that precisely identify the exact pages containing ground truth evidence. Given that documents in MMDocIR contain 65.1 pages on average, pinpointing relevant pages is highly non-trivial, akin to finding a needle in haystack, which demands careful inspection and document understanding.
@@
To enhance the granularity of our benchmark, we extend our annotations to include layout-level labels, identifying specific layout elements as evidence. Compared to page annotation, layout-level labeling is significantly more complex and labor-intensive.
  • Layout Detection. We begin by utilizing MinerU Wang et al. (2024) to automatically parse all documents and detect all layouts (e.g., layout type and bounding boxes).
  • Evidence Identification. We identify the layouts that contain necessary answer evidence. In case where MinerU fails to detect evidentiary element, we manually annotate the the bounding boxes, accounting for 7% of the total layout-level labels.
@@ 3.3 Quality Control
To ensure annotation quality and reliability in MMDocIR, we have adopted a rigorous 3-stage quality control process. We split questions into two parts. Each group is responsible for annotating approximately 1,000 questions, with an overlap of 400 questions serving the need for cross-validation.
@@ 3.4 Multimodal content as OCR/VLM-text
For layouts, the average word length and distribution of OCR-text and VLM-text of MMDocIR are shown in Figure 3(a) and 3(b). Notably, the length of VLM-text is 1.5 and 3.8 times of OCR-text for table (with more structured numbers) and figure (mostly with visual elaboration), respectively.
@@ 3.5 Statistics and Analysis
Document Analysis. As shown in Table 3, MMDocIR evaluation set includes 313 long documents, averaging 65.1 pages, categorized into 10 domains. Different domains feature distinct multimodal distribution. The overall modality distribution is as follows: text (60.4%), image (18.8%), table (16.7%), and others (4.1%), with fine-grained distribution shown in Figure 4(a).
Question and Annotation Analysis. MMDocIR includes 1,658 questions, and 2,107 page and 2,638 layout labels. The evidence spans 4 modalities: text (44.7%), image (21.7%), table (37.4%), and layout/meta (11.5%). Notably, MMDocIR presents several challenges: 254 questions require cross-modal understanding, 313 questions require evidence across multiple pages, and 637 questions require reasoning over multiple layouts.
@@ 4.2 Label Construction and statistics
Table 5: Main results for page-level retrieval, with the best results in boldface and second best results underlined. For clarity, we omit results using VLM-text (Refer to Table A for full results).
@@ 5 Experiment
Table 6: Main results for layout-level retrieval (Refer to Table A for full results with VLM-text and Hybrid inputs).
@@ 5.1 Evaluation Metric
The retriever scores each page or layout in the document based on its relevance to the question, and returns the top $k$ candidates with the highest scores. Recall@$k$ is defined as the proportion of ground truth page/layout evidence successfully retrieved. For layout matching, we calculate recall based on the overlaps between the bounding boxes of retrieved layouts and gold-standard layouts.
@@ 5.2 Baseline Models and Setting
  • Layout Retrieval: Text retrievers process multimodal layouts using OCR or VLM text (see Section 3.4). Visual retrievers process textual layouts using either Image input (cropped image of textual area) or Hybrid input (original text, as VLM can directly encode text).
@@ 5.3 Main Results for Page-level Retrieval
Table 4.2 presents the main results for page-level retrieval. Our key findings are as follows:
@@
  • Necessity of Token-level Embeddings: Compared to dense-level retrievers (e.g., BGE, DSE, DPR-Phi3ours), token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) achieve more advantageous results in Recall@1 and have marginal performance improvement in Recall@3/5. However, token-level embedding can incur storage costs of 10 times more than a single embedding (DSE requires 0.24GB for indexing MMDocIR while ColPali requires 10.0GB).
@@ 5.4 Main Results for Layout-level Retrieval
  • Necessity of Token-level Embeddings: For layout retrieval tasks, token-level retrievers marginally outperform dense-level retrievers, demonstrating its importance of in such task.
@@ 5.5 Text Retrieval: OCR-text vs VLM-Text
Most text retrievers based on on BERT Devlin et al. (2019), truncate input that exceed 512 tokens (approximately 380 english words). As shown in Figure 3(c), there are many pages containing more than 380 words (62.9% for OCR-page and 61.1% for VLM-page). Those pages suffer from critical information loss during page retrieval if the ground truth evidence is in the truncated part. In contrast, only a small fraction of layouts contain more than 380 tokens (3.9% for OCR-text, 4.8% for VLM-text, 0.5% for natural-text). Hence, as reflected in Table 4.2 and 5, text retriever demonstrates stronger performance on layout-level retrieval than on page-level retrieval.
{tabu}
ll|ccc
@@ 5.6 Visual Retrieval: Image vs Hybrid input
Visual retrievers tend to perform better when encoding text as images via visual encoders, rather than processing native textual input with LLM backbones. This advantage largely stems from their training setup: visual retrievers are typically optimized using text queries paired with image-based passages or documents, but are not fine-tuned directly on purely textual passages. However, encoding text as images incurs substantial computational overhead. Representing text as image tokens requires significantly more resources than native text encoding. To address this inefficiency and promote balanced retrieval capabilities, we advocate for future visual retrievers to be jointly trained or fine-tuned on both text and visual retrieval tasks. Such hybrid training would enable models to efficiently process text when appropriate, without compromising performance on visual inputs.
{tabu}
l|ll|ccc Model Store Index Search
(GB) (MM:SS) (MM:SS)
Page Processing
Text
DPR 0.06 6:53 00:02
@@
Layout Processing
Text
DPR 0.51 41:33 00:15
@@ 5.7 Efficiency Analysis
We evaluate the inference efficiency by measuring three key metrics : storage consumption, indexing time and, retrieval latency, as shown in Table  5.6. Experiments are conducted on a NVIDIA A100 40GB GPU with batch size of 4 for image and 256 for text. DPR-styled retrievers which generate single vector embeddings, demonstrates higher efficiency and lower computation across all metrics, compared to ColBERT-styled retrievers that produce token-level embeddings. Although DPR-styled retrievers slightly underperform in retrieval accuracy, their smaller embeddings size provide a significant advantage in the inference stage when storage space and inference time are concerned.

### v3

@@ Abstract
[…]
Figure 2: Area ratio of different modalities (1) in overall and (2) by domains in MMLongBench-Doc benchmark. Note that the white spaces, headers, and footers are removed from the area counting.
@@ 1 Introduction
Wiki-$S$ Natural question 3,610 TXT Wikipedia 1.0
@@
Multimodal document retrieval Hassan et al. (2013); Lee et al. (2024) aims to retrieve information from visually rich documents based on user queries. Unlike traditional document retrieval Zhang et al. (2022); Chen et al. (2023); Dong et al. (2024); Wang et al. (2023) and long-context Yang et al. (2025) QA which primarily deals with textual data, multimodal document retrieval imposes substantially greater demands on understanding multimodal elements such as images, tables, charts, and layout designs. Such elements often carry significant information that plain text fails to convey Cui et al. (2021); Sassioui et al. (2023): tables reveal structured data patterns, charts visualize trends or correlations, images offer contextual and semantic cues, etc. Combining these visual elements enriches the quality of retrieved content. Our analysis of MMLongBench-Doc benchmark Ma et al. (2024b) in Figure 2 shows that: text occupies only 52.7% of content area, while images and tables account for 29.2% and 12.8% respectively. This highlights the need for retrieval systems that effectively handle multimodal and cross modal Zhang et al. (2025); Dong et al. (2025b) information.
@@
However, as shown on Table 1, existing benchmarks exhibit several critical limitations that undermine comprehensive evaluation of multimodal retrieval systems. The key limitations include: 1. Question Quality: Many questions used in existing benchmarks are directly sourced from datasets for Visual Question Answering (VQA) tasks. Some questions often assume the input is already relevant, making it not suited for meaningful evaluation of retrieval capabilities. 2. Document Completeness and Diversity: Existing benchmarks often provide only partial documents, limiting the ability to evaluate within full document context. Additionally, the narrow range of document domains further restricts their applicability across diverse use-cases in real-world. 3. Retrieval Granularity: Most benchmarks support only page-level retrieval. Such granularity is often insufficient, as user queries frequently target specific elements, such as figures or tables, rather than entire pages.
@@
To address these gaps, we introduce MMDocIR, a multimodal document information retrieval benchmark. MMDocIR is designed for two key tasks: page-level and layout-level retrieval. (1) The page-level retrieval identifies the most relevant pages within a document to answer user query. (2) The layout-level retrieval targets the most relevant layouts. A layout is an element on the document page where the element could be a paragraph, a heading, an equation, a table, a figure, or a chart (see Appendix E.2 for more examples). Such task supports more precise and context-aware retrieval that pinpoint specific elements to address user queries. To support both tasks, we develop MMDocIR evaluation set that comprises 313 documents, each averaging 65.1 pages, along with 1,658 modified queries derived from MMLongBench-Doc and DocBench Zou et al. (2024). The queries are annotated with 2,107 page-level and 2,638 layout-level labels. The page labels are specific pages that contain the evidence needed to answer the query. [^1] While MMLongBench-Doc provided initial page labels, our meticulous review lead to corrections in 21.3% of them. The layout labels consist of precisely drawn bounding boxes around the key evidence within the identified pages. In addition, we introduce the MMDocIR training set, designed to support retriever training. It contains 73,843 questions sourced from 7 DocQA datasets. To construct this set, we manually collect 6,878 documents and apply a semi-automatic pipeline to annotate the ground truth labels.
@@
  • Dual-task Retrieval Framework: We propose a dual-task retrieval framework ($\S$ 2) that supports page-level and fine-grained layout-level multimodal document retrieval.
  • MMDocIR Benchmark: We introduce Multimodal Document Information Retrieval benchmark. The evaluation set ($\S$ 3) consists of 313 documents with expert-annotated labels for 1,658 questions. The training set ($\S$ 4) consists of 6,878 documents and labels for 73,843 questions.
  • We conduct extensive experiments and comparisons of both text and visual retrievers ($\S$ 5), demonstrating clear advantage of incorporating visual content in multimodal retrieval tasks.
@@ 2 Dual-Task Retrieval Definition
Let $\mathcal{D}$ be a document corpora consisting of document pages: $\mathcal{P}=\{p_{1},p_{2},\dots,p_{n}\}$, and layouts: $\mathcal{L}=\{l_{1},l_{2},\dots,l_{m}\}$ extracted via layout detection. [^2] A document page usually comprises about 5 to 15 layouts, depending on its complexity.. The objective is to perform document retrieval at both page-level and layout-level. Specifically, given query $Q$, the task is to retrieve the top $k$ pages and layouts most relevant to $Q$, where $k<<n$ and $k<<m$. The relevance of pages ($p$) and layouts ($l$) to $Q$ is measured by similarity scores, $\textrm{Sim}(Q,p)$ and $\textrm{Sim}(Q,l)$ respectively. The retrieval system consists of two phases: (1) an offline indexing phase, where pages and layouts from $\mathcal{P}$ and $\mathcal{L}$ are encoded into vectors, and (2) an online querying phase, in which a query $Q$ is encoded into a vector, which is then compared against the offline-indexed vectors using similarity scores $\textrm{Sim}(Q,p)$ for pages and $\textrm{Sim}(Q,l)$ for layouts.
@@ 3.2 Annotation Process
We annotate page labels that precisely identify the exact pages containing ground truth evidence. Given that documents in MMDocIR contain 65.1 pages on average, pinpointing relevant pages is highly non-trivial, akin to finding a needle in haystack, which demands careful inspection and document understanding. Our annotation process is described as follows:
@@
Through these efforts, we successfully obtain page-level labels for a total of 1,658 questions in MMDocIR.
@@
To enhance the granularity of our benchmark, we extend our annotations to include layout-level labels, identifying specific layout elements as evidence. Compared to page annotation, layout-level labeling is significantly more complex and labor-intensive. Our method of annotations is as follows:
  • Layout Detection. We begin by utilizing MinerU Wang et al. (2024) to automatically parse all documents and detect all layouts (e.g., layout type and bounding boxes). [^4] These elements are categorized into five main types: text, image, table, title, and equation. The bounding boxes are represented by coordinates of top-right and bottom-left corners.
  • Evidence Identification. We identify the layouts that contain necessary answer evidence. In case where MinerU fails to detect evidentiary element, we manually annotate the bounding boxes, accounting for 7% of the total layout-level labels.
Ultimately, this meticulous process leads to the annotation of 2,638 layout labels in MMDocIR.
@@ 3.3 Quality Control
To ensure annotation quality and reliability in MMDocIR, we have adopted a rigorous 3-stage quality control process. We split questions into two parts. Each group is responsible for annotating approximately 1,000 questions, with an overlap of 400 questions serving the need for cross-validation. The process for quality control works as follows:
@@ 3.4 Multimodal content as OCR/VLM-text
For layouts, the average word length and distribution of OCR-text and VLM-text of MMDocIR are shown in Figure 3(a) and 3(b). Notably, the length of VLM-text is 1.5 and 3.8 times of OCR-text for table (with more structured numbers) and figure (mostly with visual elaboration), respectively. [^5] Tables typically contain structured numbers, which can be well recorded by text. Figures contain mostly visual elaboration, making it hard to extract raw text.
@@ 3.5 Statistics and Analysis
Document Analysis. As shown in Table 3, MMDocIR evaluation set includes 313 long documents, averaging 65.1 pages, categorized into 10 domains. Different domains feature distinct multimodal distribution. For instance, reports, tutorials, workshops, and brochures predominantly contain images, whereas financial and industry documents are table-rich. In contrast, government and legal documents primarily comprise text. The overall modality distribution is as follows: text (60.4%), image (18.8%), table (16.7%), and others (4.1%), with fine-grained distribution shown in Figure 4(a).
Question and Annotation Analysis. MMDocIR includes 1,658 questions, and 2,107 page and 2,638 layout labels. The evidence spans 4 modalities: text (44.7%), image (21.7%), table (37.4%), and layout/meta [^6] The “layout/meta” refers to evidence related to layout information and meta-data statistics. (11.5%). Notably, MMDocIR presents several challenges: 254 questions require cross-modal understanding, 313 questions require evidence across multiple pages, and 637 questions require reasoning over multiple layouts.
@@ 4.2 Label Construction and statistics
Table 5: Main results for page-level retrieval, with the best results in boldface and second best results underlined. For clarity, we omit results using VLM-text (Refer to Table 11 for full results).
@@ 5 Experiment
Table 6: Main results for layout-level retrieval (Refer to Table 12 for full results with VLM-text and Hybrid inputs).
@@ 5.1 Evaluation Metric
The retriever scores each page or layout in the document based on its relevance to the question, and returns the top $k$ candidates with the highest scores. Recall@k@k is defined as the proportion of ground truth page/layout evidence successfully retrieved. For page matching, the recall is straightforwardly computed with page indices. For layout matching, we calculate recall based on the overlaps between the bounding boxes of retrieved layouts and gold-standard layouts. Unlike page retrieval, where boundaries are unambiguous, layout detection tools can produce differing bounding boxes for the same content. Our ground-truth layouts include both MinerU outputs and manual annotations for cases where MinerU misses elements. During our evaluation, retrievers are provided with MinerU predicted layouts (e.g., bboxes and types), but some ground truth bboxes cannot be exactly matched to them. Therefore, simple binary classifications (matched or not matched) are insufficient. Overlap-based recall offers a nuanced and realistic evaluation, especially where perfect alignment is not guaranteed.
@@ 5.2 Baseline Models and Setting
  • Layout Retrieval: Text retrievers process multimodal layouts using OCR or VLM text (see Section 3.4). Visual retrievers [^7] Most visual retrievers are not explicitly trained on text query-doc pairs, this setup constitutes out-of-domain data. process textual layouts using either Image input (cropped image of textual area) or Hybrid input (original text, as VLM can directly encode text).
@@ 5.3 Main Results for Page-level Retrieval
Table 5 presents the main results for page-level retrieval. Our key findings are as follows:
@@
  • Effectiveness of VLM-Text: Although VLM-text approaches underperform visual retrievers, they achieve much better performance than the OCR-text methods. This indicates benefits of using GPT-4o to preserve visual cues in text.
  • Effect of Token-level Embeddings: Compared to dense-level retrievers (e.g., BGE, DSE, DPR-Phi3ours), token-level retrievers (e.g., ColBERT, ColPali, Col-Phi3ours) achieve more advantageous results in Recall@1 and have marginal performance improvement in Recall@3/5. However, token-level embedding can incur storage costs of 10 times more than a single embedding (DSE requires 0.24GB for indexing MMDocIR while ColPali requires 10.0GB).
@@ 5.4 Main Results for Layout-level Retrieval
  • Comparison of Hybrid vs. Pure Image Sequences: Visual retrievers relying on hybrid image-text sequences generally perform less effectively than those utilizing pure image sequences. This suggests that current VLMs may have stronger capabilities in modeling images than text within the multimodal framework.
  • Effect of Token-level Embeddings: For layout retrieval tasks, token-level retrievers marginally outperform dense-level retrievers, demonstrating its importance of such task.
@@ 5.5 Text Retrieval: OCR-text vs VLM-Text
Most text retrievers based on on BERT Devlin et al. (2019), truncate input that exceed 512 tokens (approximately 380 english words). As shown in Figure 3(c), there are many pages containing more than 380 words (62.9% for OCR-page and 61.1% for VLM-page). Those pages suffer from critical information loss during page retrieval if the ground truth evidence is in the truncated part. In contrast, only a small fraction of layouts contain more than 380 tokens (3.9% for OCR-text, 4.8% for VLM-text, 0.5% for natural-text). Hence, as reflected in Table 5 and 6, text retriever demonstrates stronger performance on layout-level retrieval than on page-level retrieval.
@@ 5.6 Visual Retrieval: Image vs Hybrid input
Visual retrievers tend to perform better when encoding text as images via visual encoders, rather than processing native textual input with LLM backbones. This advantage largely stems from their training setup: visual retrievers are typically optimized using text queries paired with image-based passages or documents, but are not fine-tuned directly on purely textual passages. However, encoding text as images incurs substantial computational overhead. Representing text as image tokens requires significantly more resources than native text encoding. To address this inefficiency and promote balanced retrieval capabilities Dumitru et al. (2025); Liang et al. (2025), we advocate for future visual retrievers to be jointly trained on both text and visual retrieval tasks using SFT or RL Duong et al. (2025). Such hybrid training would enable models to efficiently process text when appropriate, without compromising performance on visual inputs.
5.7 Cascade Retrieval
As shown in Table 6, directly performing layout retrieval can be challenging. Hence, we propose alternative methods, by perform page-retrieval first, subsequently followed by layout-retrieval within the retrieved page. We term such retrieval to be cascade retrieval. Note that the page retrieval is not perfect, such error can propagate to layout retrieval and affect the final results.
page(1st) layout(2nd) Top1 Top5 Top10
BGE: direct layout 29.0 53.2 62.4
BGE BGE 24.3 49.0 58.6
E5: direct layout 26.4 51.8 61.2
E5 E5 22.2 47.8 58.2
ColBERT: direct layout 31.4 56.0 63.7
ColBERT ColBERT 28.5 53.0 61.3
ColPali: direct layout 32.5 54.3 63.2
ColPali ColPali 32.7 54.5 63.2
ColPali ColBERT 31.8 57.0 64.2
DSE: direct layout 29.1 51.9 59.9
DSE DSE 29.6 54.0 62.0
DSE ColBERT 29.0 56.4 64.4
DPR-Phi3: direct layout 30.2 53.9 62.8
DPR-Phi3 DPR-Phi3 31.1 54.2 61.7
DPR-Phi3 ColBERT 30.6 56.6 64.5
Col-Phi3: direct layout 31.6 54.5 63.3
Col-Phi3 Col-Phi3 33.3 58.6 63.7
Col-Phi3 ColBERT 35.3 58.8 65.4
Table 9: Comparison of 1-stage vs 2-stage approaches across different models
In this setting, we retrieve top-$k$ pages first, then rerank all layouts belonging to retrieved $k$ pages, and get top-$k$ layouts. The cascade retrieval results are shown in Table 9. We can observe that method with high page retrieval recall can significantly improve layout retrieval in the reranking paradigm.
Model Store Index Search
(GB) ($M$:$S$) ($M$:$S$)
Page Processing Text DPR 0.06 6:53 00:02
@@ 5.7 Cascade Retrieval
Layout Processing Text DPR 0.51 41:33 00:15
@@ 5.8 Efficiency Analysis
We evaluate the inference efficiency by measuring three key metrics : storage consumption, indexing time and, retrieval latency, as shown in Table  10. Experiments are conducted with batch size of 4 for image and 256 for text. DPR-styled retrievers which generate single vector embeddings, demonstrates higher efficiency and lower computation across all metrics, compared to ColBERT-styled retrievers that produce token-level embeddings. Although DPR-styled retrievers slightly underperform in retrieval accuracy, their smaller embeddings size provide a significant advantage in the inference stage when storage space and inference time are concerned.